# Learn2Reg 2024: New Benchmark Datasets Driving Progress on New Challenges

### v2

@@ 4.1.2 LUMIR
zhuoyuanw210 employed a dual-stream pyramid encoder paired with a GPU-accelerated motion decomposition Transformer (ModeTv2) decoder (Wang et al., 2024). The dual-stream architecture extracted hierarchical features from both moving and fixed images, generating deformations from coarse to fine across multiple scale levels.
Again, we refer to Chen et al. (2025) for comprehensive technical details and architectural specifications for all participating methods.
@@
The results demonstrate strong performance across both learning-based and classical approaches, with honkamj achieving the highest overall rank (0.814), followed by next-gen-nn (0.781) and MadeForLife (0.737) completing the top three. Notably, the top-performing methods achieved remarkably similar performance across all metrics, with the leading submissions showing narrow ranges in Target Registration Error (honkamj: 3.09, MadeForLife: 3.07, next-gen-nn: 3.12), Dice scores (0.773-0.785), and Hausdorff distances (3.04-3.29). This tight clustering of results suggests that multiple architectural strategies can achieve comparable accuracy for unsupervised brain MRI registration when properly optimized. The LUMIR challenge highlighted the significant impact of dataset scale on method performance. With over 4,000 brain MRI scans available for training, a clear trend emerged favoring deep learning approaches over classical optimization-based methods. Non-deep learning methods including VROC (rank: 0.351), ANTsSyN (rank: 0.265), and deedsBCV (rank: 0.423) consistently ranked in the lower half of submissions, contrasting with other Learn2Reg 2024 tasks where classical methods remained competitive with learning-based approaches. Foundation model approaches showed mixed results. While uniGradICONiso (rank: 0.668) with instance optimization performed reasonably well, the base uniGradICON model (rank: 0.402) ranked lower, emphasizing the importance of task-specific adaptation strategies. LoRA-FT’s parameter-efficient adaptation approach (rank: 0.384) demonstrated the potential for foundation model fine-tuning, though further optimization may be needed to compete with top-performing specialized architectures. The challenge successfully demonstrated the viability of unsupervised approaches for large-scale brain MRI registration, with all methods substantially outperforming the unregistered baseline (initial TRE: 4.38, DSC: 0.555). The removal of label supervision during training encouraged the development of more generalizable models while maintaining competitive registration accuracy, as evidenced by the consistently strong Dice scores (>0.73>0.73) among the top ten methods. Chen et al. (2025) provides deeper insights into foundational model performance, zero-shot generalization capabilities, and out-of-domain evaluation, further exploring the pathway toward robust foundational registration models for brain MRI.
@@ 4.1.3 COMULISglobe
One of the two baseline methods in the test phase was the affine variant of greedy, a fast, CPU-based deformable image registration algorithm developed by Paul Yushkevich at the Penn Image Computing and Science Lab. This variant focuses on efficient affine registration using a non-symmetric, gradient descent-based optimization, prioritizing speed over symmetry. While greedy shares conceptual roots with ANTs’ SyN (Avants et al., 2008), its streamlined design and optimized computations enable rapid alignment suitable for large-scale tasks. Serving as a classical baseline, greedy provides robust reference for comparing the effectiveness of more complex registration methods in the COMULISglobe challenge task [^10] https://github.com/hansenlasse/L2R-COMULIS-baseline-greedy.
@@
The second baseline method, globalign, performs multimodal image alignment by globally maximizing mutual information (MI) using a novel cross-mutual information function (CMIF) computed efficiently in the frequency domain (Öfverstedt et al., 2022). Designed primarily for low-degree-of-freedom transformations such as rigid alignment, globalign has demonstrated robust and fast performance across various multimodal biomedical image datasets, including histological images. Its algorithmic efficiency enables rapid estimation of global optima, contrasting with typical local MI maximization approaches. For this challenge, the default configuration—featuring a faster but less comprehensive search—was selected as the baseline. Adjustments to the hyperparameters may lead to improved results. The method’s open-source implementation is publicly available [^11] https://github.com/MIDA-group/globalign.
@@
One of the three submissions, VROC, is a classical variational image registration framework that supports flexible choices of similarity measures, regularization strategies, and transformation models (Werner et al., 2014; Ehrhardt et al., 2015). For this challenge task of multimodal registration of BF and SHG images, VROC combines handcrafted preprocessing with affine alignment and multi-resolution variational refinement. A custom-trained U-Net is used to synthesize SHG-like representations from BF inputs, reducing the modality gap. During inference, VROC performs extensive hyperparameter tuning, including affine step sizes, smoothness parameters, and optimization schedules, optionally preceded by a pre-registration step. The optimization uses normalized cross-correlation (NCC) as similarity metric, and incorporates a Demons-based strategy (Thirion, 1998) for deformable refinement. This highly engineered pipeline enables robust adaptation to the modality-specific challenges of SHG/BF registration. The general VROC framework is publicly available on GitHub [^12] https://github.com/IPMI-ICNS-UKE/vroc.
@@
next-gen-nn, employs a hybrid pipeline that combines learning-based feature extraction with classical optimization. The feature backbone, adapted from COMIR (Pielawski et al., 2020), is pretrained to extract modality-invariant representations, facilitating consistent feature alignment across SHG and BF inputs. Prealignment is performed using XFeat (Potje et al., 2024), a sparse feature matcher optimized for multimodal pairs. During training, the mutual information between image pairs is used both as the objective function and as a criterion for selecting optimal deformation fields. Fine-tuning is performed using a similar design approach to ConvexAdam (Siebert et al., 2024), a recent automatic registration framework.
@@
The lWM method combines classical feature matching with instance-specific deformable registration, forming a hybrid pipeline (Wodzinski and Müller, 2024; Wodzinski et al., 2024). The process begins with modality normalization to minimize appearance differences between BF and SHG images. Sparse keypoint correspondences are then extracted using the SuperPoint detector (DeTone et al., 2018) and matched via the SuperGlue (Sarlin et al., 2020) algorithm, which are both considered high-capacity learning-based matching models. To compensate for their lack of rotation and scale invariance, the method performs exhaustive matching across a range of initial rotations and scales, selecting the transformation with the highest number of valid matches. This provides a robust affine initialization. The final stage involves dense, multi-resolution deformable registration via instance optimization rather than deep learning. This choice was motivated by the limited availability of training data and the lack of pre-trained foundation models for microscopy registration, favoring a flexible, optimization-based approach with strong generalization properties.
@@
Figure 5: Results for the COMULISglobe SHG/BF task. *The mean inter-rater error was computed on a random subset of test cases. While thus not directly comparable to the evaluated methods, it provides a useful reference for the level of alignment that may be achievable.
@@
This division was also reflected in an initially introduced second COMULISglobe task—3D-CLEM registration—based on data from (Krentzel et al., 2023) which was ultimately dropped due to its high complexity. It involved automatic registration of multimodal 3D microscopy data, specifically aligning high-resolution electron microscopy (FIB-SEM and SBF-SEM) with super-resolution fluorescence light microscopy. The task benefited from domain knowledge and presented significant challenges related to large volumes and sparse correspondences, making it less accessible to the broader registration community. We assume it would have been of great interest to microscopy specialists, a group that was underrepresented in the challenge.
@@ 4.1.4 NLST
The NLST task presented a highly realistic and challenging benchmark for thoracic CT registration, incorporating both lung cancer screening and radiotherapy cases. The new and extended evaluation from 2023 benefited from a substantially enhanced dataset, with over 100 landmarks per case annotated independently by two medical experts, offering a rigorous basis for assessment. The top-performing method, LapIRNv2 (Mok and Chung, 2021), achieved the best overall rank (0.800), closely followed by the organizer-provided baseline (corrfield (Heinrich et al., 2015)) (rank: 0.670) and Birmingham (Jia et al., 2022, 2023) (rank: 0.553). These methods demonstrated low target registration errors (TRE: 1.425–1.63 mm) and excellent robustness on difficult cases (TRE30: 1.923–2.200 mm). While deep learning approaches dominated the top ranks, classical keypoint-based strategies remained competitive. Notably, the lightweight learning-based MONAI baseline [^13] https://github.com/Project-MONAI/tutorials/blob/main/3d_registration/learn2reg_nlst_paired_lung_ct.ipynb also performed well (rank: 0.452), highlighting the potential of efficient architectures. While teams such as UKE and UCLA reported higher TREs, all submissions significantly outperformed the unregistered baseline (initial TRE: 10.40 mm). Importantly, lung nodule TRE, a clinically critical metric, was included in the final ranking and aligned well with landmark-based scores (here the Top-4 had roughly 50% lower errors than MONAI and UCLA with TREnodules: 1.135–1.434 mm), supporting its continued use in longitudinal studies. All leading methods executed within clinically acceptable runtimes of several seconds, and the task’s realistic anatomical variability underlined the importance of robustness and generalizability in clinical deployment.
@@ 4.2 Learn2Reg 2024
The 2024 challenge also highlighted the absence of a standardized, robust registration framework analogous to nnUNet in segmentation. ConvexAdam (Siebert et al., 2024) represents a step toward this goal and was adopted by next-gen-nn in several tasks, highlighting promising research avenues focused on adaptable and reusable registration frameworks.
Baseline methods, particularly classical optimization-based tools, continue to perform strongly in some tasks. For instance, NiftyReg (Modat et al., 2014) achieved the best results on ReMIND2Reg, while Greedy (Yushkevich, 2020) shared second place with next-gen-nn on the COMULISglobe SHG/BF task. The LUMIR challenge was an exception, where classical baselines such as ANTs SyN were outperformed by deep learning-based methods. Notably, LUMIR had the largest dataset of all three tasks, with over 4,000 scans, which likely highlights the connection between dataset size and the success of deep learning approaches. Overall, while deep learning approaches demonstrate strong performance and continue to advance the field, the results suggest a degree of stagnation in optimization-based image registration and deep learning methods suited for smaller datasets, emphasizing the need for novel approaches that can generalize robustly across diverse clinical scenarios.
@@ 5 Conclusion and Outlook
We thank Kevin Eliceiri, Bin Li, and Adib Keikhosravi for providing the publicly available multimodal biomedical dataset (Eliceiri et al., 2021), which was used for the COMULISglobe SHG-BF registration task. Their contribution was instrumental in enabling the design and implementation of this multimodal image registration challenge.

### v3

@@ 4.1.2 LUMIR
zhuoyuanw210 employed a dual-stream pyramid encoder paired with a GPU-accelerated motion decomposition Transformer (ModeTv2) decoder (wang2024modetv2). The dual-stream architecture extracted hierarchical features from both moving and fixed images, generating deformations from coarse to fine across multiple scale levels.
@@
The results demonstrate strong performance across both learning-based and classical approaches, with honkamj achieving the highest overall rank (0.814), followed by next-gen-nn (0.781) and MadeForLife (0.737) completing the top three. Notably, the top-performing methods achieved remarkably similar performance across all metrics, with the leading submissions showing narrow ranges in Target Registration Error (honkamj: 3.09, MadeForLife: 3.07, next-gen-nn: 3.12), Dice scores (0.773-0.785), and Hausdorff distances (3.04-3.29). This tight clustering of results suggests that multiple architectural strategies can achieve comparable accuracy for unsupervised brain MRI registration when properly optimized. The LUMIR challenge highlighted the significant impact of dataset scale on method performance. With over 4,000 brain MRI scans available for training, a clear trend emerged favoring deep learning approaches over classical optimization-based methods. Non-deep learning methods including VROC (rank: 0.351), ANTsSyN (rank: 0.265), and deedsBCV (rank: 0.423) consistently ranked in the lower half of submissions, contrasting with other Learn2Reg 2024 tasks where classical methods remained competitive with learning-based approaches. Foundation model approaches showed mixed results. While uniGradICONiso (rank: 0.668) with instance optimization performed reasonably well, the base uniGradICON model (rank: 0.402) ranked lower, emphasizing the importance of task-specific adaptation strategies. LoRA-FT’s parameter-efficient adaptation approach (rank: 0.384) demonstrated the potential for foundation model fine-tuning, though further optimization may be needed to compete with top-performing specialized architectures. The challenge successfully demonstrated the viability of unsupervised approaches for large-scale brain MRI registration, with all methods substantially outperforming the unregistered baseline (initial TRE: 4.38, DSC: 0.555). The removal of label supervision during training encouraged the development of more generalizable models while maintaining competitive registration accuracy, as evidenced by the consistently strong Dice scores (>0.73>0.73) among the top ten methods.
@@ 4.1.3 COMULISglobe
One of the two baseline methods in the test phase was the affine variant of greedy, a fast, CPU-based deformable image registration algorithm developed by Paul Yushkevich at the Penn Image Computing and Science Lab. This variant focuses on efficient affine registration using a non-symmetric, gradient descent-based optimization, prioritizing speed over symmetry. While greedy shares conceptual roots with ANTs’ SyN (avants2008symmetric), its streamlined design and optimized computations enable rapid alignment suitable for large-scale tasks. Serving as a classical baseline, greedy provides robust reference for comparing the effectiveness of more complex registration methods in the COMULISglobe challenge task [^10] https://github.com/hansenlasse/L2R-COMULIS-baseline-greedy.
@@
The second baseline method, globalign, performs multimodal image alignment by globally maximizing mutual information (MI) using a novel cross-mutual information function (CMIF) computed efficiently in the frequency domain (ofverstedt2022fast). Designed primarily for low-degree-of-freedom transformations such as rigid alignment, globalign has demonstrated robust and fast performance across various multimodal biomedical image datasets, including histological images. Its algorithmic efficiency enables rapid estimation of global optima, contrasting with typical local MI maximization approaches. For this challenge, the default configuration—featuring a faster but less comprehensive search—was selected as the baseline. Adjustments to the hyperparameters may lead to improved results. The method’s open-source implementation is publicly available [^11] https://github.com/MIDA-group/globalign.
@@
One of the three submissions, VROC, is a classical variational image registration framework that supports flexible choices of similarity measures, regularization strategies, and transformation models (werner2014estimation; ehrhardt2015variational). For this challenge task of multimodal registration of BF and SHG images, VROC combines handcrafted preprocessing with affine alignment and multi-resolution variational refinement. A custom-trained U-Net is used to synthesize SHG-like representations from BF inputs, reducing the modality gap. During inference, VROC performs extensive hyperparameter tuning, including affine step sizes, smoothness parameters, and optimization schedules, optionally preceded by a pre-registration step. The optimization uses normalized cross-correlation (NCC) as similarity metric, and incorporates a Demons-based strategy (thirion1998image) for deformable refinement. This highly engineered pipeline enables robust adaptation to the modality-specific challenges of SHG/BF registration. The general VROC framework is publicly available on GitHub [^12] https://github.com/IPMI-ICNS-UKE/vroc.
@@
next-gen-nn, employs a hybrid pipeline that combines learning-based feature extraction with classical optimization. The feature backbone, adapted from COMIR (pielawski2020comir), is pretrained to extract modality-invariant representations, facilitating consistent feature alignment across SHG and BF inputs. Prealignment is performed using XFeat (potje2024xfeat), a sparse feature matcher optimized for multimodal pairs. During training, the mutual information between image pairs is used both as the objective function and as a criterion for selecting optimal deformation fields. Fine-tuning is performed using a similar design approach to ConvexAdam (siebert2024convexadam), a recent automatic registration framework.
@@
The lWM method combines classical feature matching with instance-specific deformable registration, forming a hybrid pipeline (wodzinski2024automatic; wodzinski2024deeperhistreg). The process begins with modality normalization to minimize appearance differences between BF and SHG images. Sparse keypoint correspondences are then extracted using the SuperPoint detector (detone2018superpoint) and matched via the SuperGlue (sarlin2020superglue) algorithm, which are both considered high-capacity learning-based matching models. To compensate for their lack of rotation and scale invariance, the method performs exhaustive matching across a range of initial rotations and scales, selecting the transformation with the highest number of valid matches. This provides a robust affine initialization. The final stage involves dense, multi-resolution deformable registration via instance optimization rather than deep learning. This choice was motivated by the limited availability of training data and the lack of pre-trained foundation models for microscopy registration, favoring a flexible, optimization-based approach with strong generalization properties.
@@
Figure 5: Results for the COMULISglobe SHG/BF task. *The mean inter-rater error was computed on a random subset of test cases. While thus not directly comparable to the evaluated methods, it provides a useful reference for the level of alignment that may be achievable. * Baseline method provided by organizers.
@@
This division was also reflected in an initially introduced second COMULISglobe task—3D-CLEM registration—based on data from (krentzel2023clem) which was ultimately dropped due to its high complexity. It involved automatic registration of multimodal 3D microscopy data, specifically aligning high-resolution electron microscopy (FIB-SEM and SBF-SEM) with super-resolution fluorescence light microscopy. The task benefited from domain knowledge and presented significant challenges related to large volumes and sparse correspondences, making it less accessible to the broader registration community. We assume it would have been of great interest to microscopy specialists, a group that was underrepresented in the challenge.
@@ 4.1.4 NLST
The NLST task presented a highly realistic and challenging benchmark for thoracic CT registration, incorporating both lung cancer screening and radiotherapy cases. The new and extended evaluation from 2023 benefited from a substantially enhanced dataset, with over 100 landmarks per case annotated independently by two medical experts, offering a rigorous basis for assessment. The top-performing method, LapIRNv2 (mok2021conditional), achieved the best overall rank (0.800), closely followed by the organizer-provided baseline (corrfield (heinrich2015estimating)) (rank: 0.670) and Birmingham (10.1007/978-3-031-21014-3_16; jia2023fourier) (rank: 0.553). These methods demonstrated low target registration errors (TRE: 1.425–1.63 mm) and excellent robustness on difficult cases (TRE30: 1.923–2.200 mm). While deep learning approaches dominated the top ranks, classical keypoint-based strategies remained competitive. Notably, the lightweight learning-based MONAI baseline [^13] https://github.com/Project-MONAI/tutorials/blob/main/3d_registration/learn2reg_nlst_paired_lung_ct.ipynb also performed well (rank: 0.452), highlighting the potential of efficient architectures. While teams such as UKE and UCLA reported higher TREs, all submissions significantly outperformed the unregistered baseline (initial TRE: 10.40 mm). Importantly, lung nodule TRE, a clinically critical metric, was included in the final ranking and aligned well with landmark-based scores (here the Top-4 had roughly 50% lower errors than MONAI and UCLA with TREnodules: 1.135–1.434 mm), supporting its continued use in longitudinal studies. All leading methods executed within clinically acceptable runtimes of several seconds, and the task’s realistic anatomical variability underlined the importance of robustness and generalizability in clinical deployment.
@@ 4.2 Learn2Reg 2024
The 2024 challenge also highlighted the absence of a standardized, robust registration framework analogous to nnUNet in segmentation. ConvexAdam (siebert2024convexadam) represents a step toward this goal and was adopted by next-gen-nn in several tasks, highlighting promising research avenues focused on adaptable and reusable registration frameworks.
Baseline methods, particularly classical optimization-based tools, continue to perform strongly in some tasks. For instance, NiftyReg (modat2014global) achieved the best results on ReMIND2Reg, while Greedy (greedy) shared second place with next-gen-nn on the COMULISglobe SHG/BF task. The LUMIR challenge was an exception, where classical baselines such as ANTs SyN were outperformed by deep learning-based methods. Notably, LUMIR had the largest dataset of all three tasks, with over 4,000 scans, which likely highlights the connection between dataset size and the success of deep learning approaches. Overall, while deep learning approaches demonstrate strong performance and continue to advance the field, the results suggest a degree of stagnation in optimization-based image registration and deep learning methods suited for smaller datasets, emphasizing the need for novel approaches that can generalize robustly across diverse clinical scenarios.
@@ 5 Conclusion and Outlook
We thank Kevin Eliceiri, Bin Li, and Adib Keikhosravi for providing the publicly available multimodal biomedical dataset (eliceiri2021dataset), which was used for the COMULISglobe SHG-BF registration task. Their contribution was instrumental in enabling the design and implementation of this multimodal image registration challenge.